OpenStudy (kingswank):

MEDALS WILL BE GIVEN!!! Use the net to find the surface area of the cylinder. R=5yd H=7yd.

OpenStudy (perl):

we need to see the net

OpenStudy (tkhunny):

@kingswank Please consider the terms of service and code of conduct. Stop posting problems with no work shown. You must show YOUR work. If you really have nothing to offer, you are in the wrong class and we cannot help you, here. SHOW YOUR WORK!

OpenStudy (anonymous):

Surface Area of Cylinders SA = 2πr2 + 2πrh General Cylinder General Cylinder A cylinder has a total of three surfaces: a top, bottom, and middle. The top and bottom, which are circles, are easy to visualize. The area of a circle is πr2. So, the area of two circles would be πr2 + πr2 = 2πr2. The third surface, the lateral surface area, is less easy to visualize for the purposes of calculating its area, especially since it does not appear to be in a shape that fits a known area like a triangle or parallelogram. The surface being referred to is the curved wall of the cylinder. Lateral Area of a Cylinder We can manipulate the curved wall of a cylinder to produce a recognizable shape. Imagine starting with a cylinder, like a typical can, and then cutting the can up its wall. From the bottom to the top of the can, a cut is made. See the dashed segment in step one of the diagram to the right. Now, from the newly cut edge, the cylinder wall will be spread open. Almost like opening double doors to a fancy home, the walls spread apart. See step two of the same diagram. When the cylinder's wall is completely open, it takes the form a very recognizable, basic shape. It is a rectangle. Fortunately, the area of a rectangle is easy to calculate. But we need to go back to the original cylinder for a moment before we calculate the rectangle's area. The top of the original cylinder is a circle. We know that the distance around a circle is called its circumference, C = 2πr. The circumference has been marked in red. When the cylinder wall is completely open, we see that the circumference of the circle becomes the length of the final rectangle. The dimensions of the rectangle are the circumference, C = 2πr, and the height of the cylinder, h. So, the area of the rectangle is A = l x w = C x h = 2πr x h = 2πrh. Finally, if we put all the three surfaces together, we can arrive at the formula for the complete surface area of a cylinder. The top, the bottom, and the walls of the cylinder (the lateral surface area) determine the total surface area. Those values are πr2 + πr2 + 2πrh = 2πr2 + 2πrh. Sometimes this expression is written as 2πr(r + h) in certain geometry texts. [Can you see how 2πr2 + 2πrh = 2πr(r + h) by the distributive property?] Example 1: If r = 6 ft and h = 4 ft, then the surface area would be... SA = 2πr2 + 2πrh SA = 2(3.14159)(6 ft)2 + 2(3.14159)(6 ft)(4 ft) SA = 2(3.14159)(36 ft2) + 2(3.14159)(24 ft2) SA = 226.19448 ft2 + 150.79632 ft2 SA = 376.9908 ft2 SA = 377.0 ft2 (rounding to the nearest tenth) Example 2: If r = 5 cm and h = 2 cm, then the surface area would be... SA = 2πr2 + 2πrh SA = 2(3.14159)(5 cm)2 + 2(3.14159)(5 cm)(2 cm) SA = 2(3.14159)(25 cm2) + 2(3.14159)(10 cm2) SA = 157.0795 cm2 + 62.8318 cm2 SA = 219.9113 cm2 SA = 219.9 cm2 (rounding to the nearest tenth)

OpenStudy (kingswank):